OpenStudy (briana.img):

Solve the System y= 4x + 12 y= -2x +6

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

4x+12=-2x+6 +2x +2x ------------ 6x+12=6 -12 -12 ----------- 6x=6 -- -- 6 6 x=1

OpenStudy (briana.img):

@ohlookidontunderstandanotherthing i need an ordered pair

OpenStudy (anonymous):

Ive no clue then. Sorry! I tried!

OpenStudy (briana.img):

@ohlookidontunderstandanotherthing it's okay!!

OpenStudy (here_to_help15):

ill try

OpenStudy (here_to_help15):

x=-1,y=8

OpenStudy (here_to_help15):

do u need steps?

OpenStudy (briana.img):

@Here_to_Help15 please!

OpenStudy (here_to_help15):

here umm go on Google and type system calculator click on the 3rd 1 and write your problem there it will solve it and offer the steps click it and there you go :)

OpenStudy (here_to_help15):

i didnt feel like writing the steps out again sorry :(:)

OpenStudy (here_to_help15):

and are you taking algebra b?

OpenStudy (briana.img):

@Here_to_Help15 yeah almost done with the course

OpenStudy (here_to_help15):

me to :P

OpenStudy (here_to_help15):

well i finished it :P

OpenStudy (here_to_help15):

Are u taking algebra early or on time?

OpenStudy (briana.img):

@Here_to_Help15 retaking the class

OpenStudy (here_to_help15):

oh so then what grade r u in?

OpenStudy (briana.img):

@Here_to_Help15 going into my junior year

OpenStudy (here_to_help15):

so ur going to 9th?

OpenStudy (briana.img):

@Here_to_Help15 11th

OpenStudy (here_to_help15):

i meant 11th hahaha

OpenStudy (here_to_help15):

im going to 10th

OpenStudy (here_to_help15):

Whats your fav hobby?

OpenStudy (briana.img):

@Here_to_Help15 make videos..

OpenStudy (here_to_help15):

oh ur just like my bros:P

OpenStudy (here_to_help15):

ummm.. sport?